Fermi National Laboratory
 

Accelerator Update for January 3, 2007 - January 5, 2007

 

Wednesday January 3

The day shift began with Operations monitoring store 5160 (52.36E30) and a Recycler stash (268.4E10, with the Antiproton Source (Pbar) stacking (38.3E10), with Linac, Booster, Main Injector (MI), and Recycler operating normally, with MiniBooNE off due to absorber work, with NuMI taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beams.

MI experts began a collimator study and Booster experts began a quadrupole damper study at 9:41 AM.

A Booster RF station (BRF6) tripped off for the first of three times in about an hour and a half on gap envelope faults. It reset each time.

At 10:32 AM, some Pbar Debuncher TWTs tripped off due to a problem with an RF station (DRF1). Experts investigated and found a bad connector on a NIM card. Normal stacking resumed at 11:36 AM.

At 12:46 PM, a Tevatron (TeV) air compressor (T:C4AIRP) alarmed. It reset.

The Recycler completed its second and last stash of antiprotons for the day shift at 1:23 PM.

Operations terminated store 5160 at 1:32 PM.

MI experts completed their collimator study at 1:41 PM.

At 1:45 PM, T:C4AIRP tripped off again. Operators investigated and then reset the compressor.

Operations put the TeV into shot setup at 1:54 PM.

At 2:41 PM, several Booster devices began having readback problems. Booster experts investigated and later resolved the problem by rebooting the Booster controls system.

Operations established store 5161 at 4:48 PM with an initial luminosity of 233.95E30.

The Recycler completed its third a final stash for the evening shift at 10:38 PM.

At 11:19 PM, BRF6 tripped off. It reset okay.

Thursday January 4

The midnight shift began with Operations monitoring store 5161 (95.4E30) and the stash (128.2E10), with Pbar stacking (37.8E10), and with NuMI taking beam.

At 1:04 AM, BRF6 tripped off. It reset okay.

At 6:03 AM, BRF7 & 8 tripped off due to high vacuum pressure (IPRF17). They reset okay.

At 6:08 AM, LRF5 tripped off due to a power amplifier (PA) crowbar and would not reset. Operations investigated and reported hearing sparking. The station came back online around 6:45 AM.

At 8:45 AM, Operations reported that LRF5 was continuing to trip off. Operators investigated and contacted Linac technicians.

Booster RF experts took BRF6 offline at 9:04 AM to make repairs. It was back online 35 minutes later.

Operations had LRF5 back on by 9:07 AM.

MI experts began barrier bucket studies at 12:21 PM.

The Neutron Therapy Facility (NTF) reported at 12:44 PM that they had finished treating their last patient for the day.

MI experts completed their studies at 3:36 PM.

Operations terminated store 5161 at 4 PM. Operations reported that the Recycler had completed 3 stashes for the day shift.

At 4:20 PM, LRF5 tripped off. It reset okay.

Operations put the TeV into shot setup at 4:23 PM.

At 5:29 PM, Operations received a report of a Linac LCW leak from KRF6. Operators discovered that the leak was rainwater coming from the ceiling. Operators and the Duty Mechanic diverted the leak away from the RF station.

Operations established store 5162 at 6:37 PM with an initial luminosity of 220.06E30.

At 7:47 PM, the Preaccelerator's H- Source chopper tripped off due to a vacuum valve closing.

At 8:29 PM, the Pbar Debuncher cooling amplifiers at AP-30 tripped off. Operations called in an expert who found a blown fuse for the bulk power supply.

The Recycler completed its second and last stash for the evening shift at 9:10 PM.

Friday January 5

The midnight shift began with Operations monitoring store 5162 (103.22E30), with the Recycler and Pbar preparing to stash, and with NuMI taking beam.

The Recycler had completed its third and last stash for the midnight shift around 5 AM.

At 5:23 AM, the NuMI beam permit tripped off due to a beam position problem. It reset okay.

At 6:35 AM, a Recycler RF cooling station (R:AR5KW1) tripped off. A Recycler expert switched to an alternate way of cooling the beam until the RF station gets repaired.

The Plans for Friday and the Weekend

The plans for today and the weekend are to stack, stash, and store. The Run Coordinator will allow MI and Booster experts to conduct studies this morning and possibly a Linac study this afternoon.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 01/05/2007    email Fermilab